10. A company has with a par value of $1,000 which pay semiannual interest of $60. The bonds are set to mature in 15 years and the interest rates on comparable bonds (YTM) is 10% per annum. What is the price of this bond?
Answer by Theo(13342)

the formula i found on https://www.wallstreetmojo.com/bond-pricing-formula/ is:

bond price = C * (1 - (1+r)^-n)/r + F/1+r)^n

C is the coupon price
r is the market interest rate per time period
n is the number of time periods
F is the par value of the bond.

in your problem:

C = 60
r = .10 per year / 2 = .05 per semi-annual period.
n = 15 years * 2 = 30 semi-annual periods.
F = 1000

in your problem:
bond price = C * (1 - (1+r)^-n)/r + F/1+r)^n becomes:
bond price = 60 * (1 - (1+.05)^-30)/.05 + 1000/(1+.05)^15 = 1153.72.

1153.72 would be the price of the bond.

i got the formula from https://www.wallstreetmojo.com/bond-pricing-formula/

i also used an online bond price calculator at https://exploringfinance.com/bond-price-calculator/ and got the same result.

the results from using that calculator can be found below:



your inputs to that calculator are:

number of years to maturity = 15%
yield or market rate = 10%
bond face value = 1000
annual coupon rate = 12%
coupons are given every semi-annual period.

since the coupon was 60 dollars every semi-annual time period, then the annual coupon was 120.
120 / 1000 = .12 = 12% annual coupon rate.